OpenStudy (anonymous):

Find the n th term and sum of n terms of the series 5 + 8 + 17 + 44 ..........

OpenStudy (anonymous):

it is not Ap lol

Parth (parthkohli):

\(a_n = a_{n - 1} + 3^{n - 1}\)

Parth (parthkohli):

Let me check if that is correct.

OpenStudy (jiteshmeghwal9):

geometric progression??

OpenStudy (jiteshmeghwal9):

mean

Parth (parthkohli):

Yep. The formula I posted is correct for the \(n\)th term.

OpenStudy (anonymous):

@ParthKohli wat did u did!!

Parth (parthkohli):

I just tried to see the difference. 8 - 5 = 3^1 17 - 8 = 3^2 44 - 17 = 3^3

OpenStudy (anonymous):

Nice..

Parth (parthkohli):

@waterineyes Website?

OpenStudy (anonymous):

so the 1 st order differrence is in gp

OpenStudy (anonymous):

then...

OpenStudy (anonymous):

Which website ?? @ParthKohli Ha ha ha..

Parth (parthkohli):

Use sigma to express the summation.

OpenStudy (jiteshmeghwal9):

i also asked one of the question like this i.e\[1^2+2^2....n^2\]n=?

Parth (parthkohli):

The answer to @jiteshmeghwal9's question in sigma notation is: \[\sum_{i = 1}^{n} {i^2} \] Do you want to know how I did that?

OpenStudy (anonymous):

= n(n+1) (2n+1)/6

Parth (parthkohli):

How did you get that? @Yahoo!

OpenStudy (anonymous):

it is the general solution

Parth (parthkohli):

I see.

OpenStudy (anonymous):

n^3 - (n(n+1))/2)^2

Parth (parthkohli):

@mukushla Can you help us?

OpenStudy (jiteshmeghwal9):

ya! @ParthKohli i will be interested to know:)

OpenStudy (anonymous):

@ParthKohli y got it already

Parth (parthkohli):

Trying to find the sum.

OpenStudy (jiteshmeghwal9):

@Yahoo! has gt the right answer to my question:)

OpenStudy (jiteshmeghwal9):

@ParthKohli the answer given in my book is matching with @Yahoo! 's answer:)

OpenStudy (jiteshmeghwal9):

i have determine a pattern a pattern of @Yahoo! 's question

OpenStudy (anonymous):

@jiteshmeghwal9 can u show that

Parth (parthkohli):

\(5 + 3 \Longrightarrow 2^{nd}\) \(5 + 3 + 3^2 \Longrightarrow 3^{rd}\) I am not sure how to represent these in sigma notation.

OpenStudy (anonymous):

\(a_n=a_{n-1}+3^{n-1}=a_{n-2}+3^{n-1}+3^{n-2}=...=a_1+3+...+3^{n-2}+3^{n-1} \)

Parth (parthkohli):

Any other way to represent that?

OpenStudy (jiteshmeghwal9):

wht @ParthKohli replied after ur reply @Yahoo!

OpenStudy (anonymous):